Attendees:

Notes:

  • Christopher Waters 
    • Update DMTN-222.
    • Sort out summit scripts.
      • Add slack channel for this discussion.
    • Electrometer ingest.
    • Design TAXICAB report format.
    • Summit access.
      • Jenkins access.  This is sorted out.
    • Filter scan to review.
      • The monochrometer should be linked with the fiber spectrograph.
    • More analysis_tools issues:
      • Sasquatch connection?
      • Summit status?
      • Filtering for detector level tasks.
  • Agnès Ferté 
    • Working on unit tests and isrMock updates.
    • Calibrations with new task.
  • Daniel Weatherill 
    • Looking into analysis_tools cleanup tickets.
    • Getting up to speed on the analysis_tools breaks.
  • Eli Rykoff 
    • RFC on renaming
      • Updated target names
    • Linearity ticket back into review.
      • Includes temperature fitting
      • Does not attempt to identify why the flat stability doesn't match these temperature values.
      • Computes sigma mad to see what the worst fit amplifiers are
    • Updated defects for December
    • New ISR Task planning for use.
    • Need to bootstrap a PTC pipeline to make the linearity and then apply the linearity.
      • Outlier rejection in PTC may not be appropriate for the linearity as well.  DM-44072
  • Yousuke Utsumi 
    • EO testing has based on the existing ISR task, not the new one.
      • Switch needs to be done now if needed.
      • Is this a problem?
    • Is there any metrics being generated now?
      • Not yet.
    • On the mountain this week.
  • Robert Lupton 
    • Visit database has not had anything added to it.
      • sdm_schemas?
    • Which "ComCam" data will we use for the rehearsal?